Use substitution to solve each system of equations. If the system does NOT have exactly one solution, state whether it has NO solutions, or INFINITELY MANY solutions.
1.) y=7-x
2x-y=8

Answer by stanbon(75887)

Use substitution to solve each system of equations. If the system does NOT have exactly one solution, state whether it has NO solutions, or INFINITELY MANY solutions.
1.)
y=7-x
2x-y=8
------
Substitute for "y" and solve for "x":
---
2x -(7-x) = 8
2x + x -7 = 8
3x = 15
x = 5
-----
Solve for "y":
y = 7-x
y = 7-5
y = 2
----------
Solution: (5,2)
=====================
Cheers,
Stan H.